OpenStudy (anonymous):

What is the length of a side of a cube with a volume of 729 m3 @rebeka

OpenStudy (anonymous):

is it 27

jagr2713 (jagr2713):

V = x^3 = 729 x^3 = 9^3 x = 9 m

OpenStudy (anonymous):

so 3*3*3

jagr2713 (jagr2713):

no 3*3*3 is 27

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

729/3=answer

OpenStudy (anonymous):

@rebeka

OpenStudy (rebeka):

it 9 like @jagr2713 said

OpenStudy (triciaal):